# Public Projects with Preferences and Predictions

### v1

@@ Abstract
In the public projects problem, a group of decisionmakers aggregate their preferences to choose one alternative. Recent work on public projects has proposed the Quadratic Transfers Mechanism (QTM) and shown asymptotic welfare guarantees in some cases. We begin by giving new non-asymptotic Price of Anarchy guarantees for the QTM.
We then incorporate an alternative philosophy toward group decisionmaking, aggregation of information about which is the best alternative. We propose a public projects mechanism based on the QTM that aggregates both preferences and predictions, modeled as forecasts of the projects’ welfare impacts. When the predictions come from a prediction market or wagering mechanism, we show the entire mechanism is robust to manipulation and give Price of Anarchy guarantees, though under strong assumptions on the mechanism’s knowledge. Our results focus primarily on the case of deciding between two alternatives, showing the Price of Anarchy tends to $1$ as natural measures of the “size” of the population grow large. In most cases, the mechanisms achieve a balanced budget as well.
@@ 1 Introduction
In the public projects problem, a group of decisionmakers is faced with selecting one of multiple alternative projects to undertake. Classically, the objective is to aggregate the preferences of the decisionmakers in order to make a choice which most benefits the group as a whole. This paradigm includes social choice theory and mechanism design.
@@
An alternative philosophy is to focus on aggregating information about which alternative is the best. For example, a committee could make a decision either by voting (preferences), or by discussing until rough consensus is reached (information). Decision markets [Othman and Sandholm, 2010] have been proposed as a group decisionmaking mechanism in the information paradigm. There, forecasts about the impact of each alternative are produced by running prediction markets – financial markets whose prices reflect predictions – and then a decision is automatically chosen based on the predicted-best alternative. We likewise focus on predictions as a source of credible, (partly) verifiable information.
@@
A public-projects problem can generally benefit from aggregation of both preferences and information. Can we design a mechanism for groups to make decisions based on both? There are only a few formal proposals in this spirit (see Section 1.2). When agents can serve as both information aggregators and preference-driven decisionmakers, they may have incentives to manipulate in both roles.
In this paper, we will propose a mechanism that aggregates both preferences and predictions to make a group decision. For quantitative results, we consider monetary mechanisms and assume quasilinear utility, seeking social welfare guarantees. Each participant $i$ has a value $v_{k}^{i}\geq 0$ for each project $k$, with $V_{k}=\sum_{i}v_{k}^{i}$. Meanwhile, project $k$ has some unknown expected external welfare impact $B_{k}$, a function of participants’ private information. We seek to elicit the preferences $\{v_{k}^{i}\}$ and information $\{B_{k}\}$ and maximize social welfare, $\arg\max_{k}(V_{k}+B_{k})$. The Synthetic Players Quadratic Transfers Mechanism with Predictions (SQUAP) begins with an information aggregation phase, such as a prediction market or a wagering mechanism, to elicit $\{B_{k}\}$. Then, it uses a “voting” stage, based on the recently-introduced Quadratic Transfers Mechanism discussed below, to elicit preferences and select an alternative. The mechanism is practical and budget-balanced. Unfortunately, it is naturally also very difficult to analyze strategically. However, by adding some assumptions on the designer’s knowledge, we can obtain a less-practical variant we call Impractical SQUAP. Then we can prove the following bound on the Price of Anarchy, or worst-case equilibrium ratio of the mechanism’s welfare to optimal-achievable welfare.
@@
In Impractical SQUAP, using either a prediction market or wagering mechanism, for two alternatives,
@@
More formally, $T=\frac{\arg\max_{k}(V_{k}+B_{k})}{\max_{i,k}v_{k}^{i}}$, the ratio of first-best welfare to the largest individual value. As the total welfare grows large relative to the maximium influence of any one person’s preference, the Price of Anarchy tends to one. This result proves that it is possible to align incentives when strategic agents are both forecasters and decisionmakers. We view the impractical mechanism’s welfare guarantee as strong evidence for the welfare of our main proposed mechanism.
1.1 Outline and Results
@@
Recent work [Eguia et al., 2019, 2023] has introduced and studied the Quadratic Transfers Mechanism (QTM) for public projects. Inspired by quadratic voting [Lalley and Weyl, 2018, Goeree and Zhang, 2017], the mechanism collects bids (“votes”) on each alternative, with cost rising quadratically, and picks an alternative according to a probabilistic “softmax” of the vote totals. By redistributing payments, the mechanism is budget-balanced.
@@
To analyze our combined aggregation-decision mechanisms, we first analyze the QTM and present Price of Anarchy guarantees (Section 2). Our approach and results build on the work of Eguia et al. [2019], which gives asymptotic welfare bounds in a model with agents drawn i.i.d. from bounded value distributions, where realized values are common knowledge. We consider any worst-case set of agents with no distributional assumptions. We analyze pure-strategy Nash equilibrium, a natural solution concept in this setting for several reasons:
  • Pure strategies are natural because the strategy space in the QTM is a convex set, $\operatorname{\mathbb{R}}^{m}$, the number of “votes” to cast on each alternative. Further, we show that for a reasonable setting of the QTM parameter, utilities are strictly concave, implying that if others play pure strategies, it is strictly suboptimal to randomize.
  • (Nash vs Bayes-Nash, I) In the QTM, computing a best-response only requires knowledge of opponents’ aggregate vote totals. The situation is reminiscent of mean field games [Caines, 2021]. We view this amount of knowledge as equally reasonable to a Bayes-Nash equilibrium assumption that participants know e.g. the distribution of the vote totals.
  • (Nash vs Bayes-Nash, II) A designer can facilitate the spread of information required to achieve Nash equilibrium, e.g. with straw poll, or a dynamic voting process that publicizes information about the vote totals during the mechanism. On the other hand, if a Bayes-Nash analysis were to rely on agents’ ignorance of the others’ realized types, it could be fragile.
Types of information.
Broadly, there are two types of information relevant to the decision: internal-relevant and external-relevant. Internal-relevant information affects decisionmakers’ own preferences. External-relevant information affects estimated “external welfare” of the decision, i.e. impacts on others. For example, consider shareholders of a company deciding between two product lines. The expected revenue of each product would be internal-relevant, as each shareholder’s utility and preference is related to the revenue. The climate impact (e.g. expected $\text{CO}_{2}$ emissions) would be external-relevant, as it does not impact each shareholder’s utility directly (for the sake of argument). We focus on modeling external-relevant information only, i.e. the external welfare impacts $\{B_{k}\}$. Our mechanism could easily be modified to elicit internal-relevant information in the first stage (e.g. with prediction markets). However, for the sake of analysis, we assume participant preferences $\{v_{k}^{i}\}$ are fixed independent of $\{B_{k}\}$.
External-relevant information is important in many public-projects or group decisionmaking contexts. If a governing board, shareholders, or co-op members are making a decision about their organization, their charter may require consideration of the impact on the organization as a whole, its employees, or the population the organization is intended to serve. This could hold particularly for a non-profit, co-op, or a Decentralized Autonomous Organization (DAO) in a blockchain-governance context [Spelliscy et al., 2024]. Such decisions often involve a conflict between what is best for the decisionmakers versus what is best for the organization’s mission. Our mechanism could help by aligning incentives and optimizing for a combination.
To move toward incorporating information, we consider in Section 3 an extension of the QTM to a case where the external welfare impact of each project is fixed and known. Our Synthetic Players QTM first collects agent votes, then “participates” with synthetic players whose values capture the external welfare impacts. Unfortunately, analysis is challenging. We consider an “impractical variant” that needs knowledge about the setting and commits to the synthetic votes first, then collects votes from the agents. In this case, we are again able to prove a strong Price of Anarchy bound for the two-alternative case.
@@
In Section 4 we define a two stage mechanism, Synthetic QUAdratic Transfers with Predictions (SQUAP). In the first stage, agents participate in an information-aggregation mechanism, such as prediction market or wagering mechanism to estimate the external welfare impacts $\{B_{k}\}$ of the projects. Then, the second stage runs the Synthetic Players QTM from Section 3 with input $\{B_{k}\}$ to produce the final decision.
To prove a Price of Anarchy bound (for the impractical variant), we take the following steps. First, we consider two commonly-studied mechanisms for eliciting predictions: prediction markets and wagering mechanisms. Second, we show that, while agents may have an incentive to misreport predictions given the decision is based on their information, deviations will be bounded and tend to zero as the size of the information-aggregation mechanism grows. Finally, we use an importance-weighting technique from the decision markets literature [Chen et al., 2011] so that forecasters’ expected payoffs are independent of the decision of the mechanism. Under some mild assumptions, we provide a Price of Anarchy guarantee, our main result, based both on the spread of the values $T$ and the size of the information-aggregation mechanism.
@@ 1.2 Related Work
There is naturally a significant amount of work on public projects from a mechanism-design perspective. We focus on the setting of this paper, which is social welfare maximization with quasilinear-utility agents. A standard solution is to use the Vickrey-Clarke-Groves mechanism (VCG) [Vickrey, 1961, Clarke, 1971, Groves, 1973] which has an equilibrium that is truthful and maximizes social welfare. However, VCG has some undesirable properties, including lack of a Price of Anarchy bound (there exist arbitrarily bad equilibria) and uncertain revenue (it may be very large or zero). Literature addressing the problem with more nuance than our basic model considers the excludable and non-excludable cases, e.g. Ohseto [2000]. The smoothness framework has been used to show a constant-factor Price of Anarchy bound for a mechanism in which participants place bids on all alternatives, the one with the highest total is selected, and everyone pays their bid on the winner [Roughgarden et al., 2017].
@@
Recent research has considered transfer-based mechanisms which use quadratic payments  [Weyl, 2013, Lalley and Weyl, 2018, Goeree and Zhang, 2017]. Lalley and Weyl [2018, 2019] propose a binary-outcome quadratic voting mechanisms and show it is asymptotically efficient. Weyl [2017] then studies this mechanism’s sensitivity to collusion, fraud, and voter mistakes when the number of agents is large. A number of other works study the behavior or empirical performance of quadratic voting [Chandar and Weyl, 2019, Casella and Sanchez, 2019, Quarfoot et al., 2017, Goeree and Zhang, 2017].
Building on Lalley and Weyl [2019], the QTM in particular is proposed by Eguia et al. [2019] for $m\geq 2$ alternatives. The authors consider a setting where participants’ values for the $m$ alternatives are drawn i.i.d. from a distribution over $[0,1]^{m}$, and realized values are common knowledge. They prove that for an undetermined choice of the mechanism’s scale parameter $c$, equilibria exist, and show that the probability of selecting the highest-welfare alternative approaches 1 as the number of agents diverges. Eguia et al. [2023] then study the same mechanism in a similar setting but under Bayes-Nash mixed-strategy equilibrium, with similar results. The part of our work on the QTM builds heavily on the approaches in Eguia et al. [2019, 2023], but our results address the nonasymptotic regime without distributional assumptions. This includes concrete bounds with small constants.
Aggregating information for decisionmaking.
Many works study mechanisms for eliciting information from experts and using it to make a decision. The most relevant mechanisms for this paper are decision markets [Hanson, 1999, Othman and Sandholm, 2010], which simultaneously operate a prediction market for each alternative $k$. (A prediction market is a financial market designed to aggregate beliefs into a consensus forecast, reflected in the prices of the financial products being traded [Hanson, 2003].) The decision market mechanism selects the alternative that is best according to the aggregated predictions (or uses a similar rule). Even without preferences over the alternatives, decision markets can have complex incentive misalignment problems, although these can be fixed by the importance-weighting technique of Chen et al. [2011].
@@
There are many works in which a single decisionmaker elicits information from one or more experts prior to making a decision, and the experts wish to influence the decision. This includes the “cheap talk” model of Crawford and Sobel [1982] as well as Bayesian Persuasion [Kamenica and Gentzkow, 2011]. Oesterheld and Conitzer [2020] considers “decision scoring rules” that incentivize truthful predictions and recommendations by an expert.
For aggregating information from a group, e.g. Gerardi et al. [2009], consider a Bayesian setting in which a decisionmaker elicits information from a number of experts and aggregates it to make a decision. To incentivize truthfulness, the decisionmaker with small probability audits an expert’s report against the others (similar to peer prediction [Miller et al., 2005]) and picks an alternative desired by the audited agent.
Such settings typically do not face the concern in this paper, where the decisionmakers are a large public group and may overlap with the forecasters.
@@
A few works are much closer to the spirit of our motivation, if technically quite different. They generally analyze voter preferences as a function of the information to be aggregated. In contrast, we model preferences as fixed with information as an orthogonal axis. In the “Wisdom-of-the-Crowd Voting Mechanism” of Schoenebeck and Tao [2021], each participant is both a voter and a holder of private information. The mechanism elicits both information and a preference in a single shot, then aggregates both to make a binary decision. The authors show that, with high probability, the mechanism selects despite strategic behavior the “majority wish” alternative that more voters would prefer if fully informed.
@@
Amanatidis et al. [2022] consider a similar problem motivated by blockchain applications. Voters have private information and participate in an approval vote among $m\geq 2$ alternatives. As with Schoenebeck and Tao [2021], information aggregation takes place within the mechanism rather than through an explicit phase prior to voting. The authors prove that the mechanism can achieve a Price of Anarchy of $\frac{1}{2}$.
@@
In Jackson and Tan [2013], a majority-vote between two alternatives takes place after a “deliberation” stage. The experts who take part in deliberation cannot also be voters (unlike in our model), and cannot misreport arbitrarily, but only choose to either reveal or hide their information in order to influence the voters. Other works in this spirit are Alonso and Câmara [2016], Schnakenberg [2015].

### v2

@@ Abstract
When making a decision as a group, there are two primary paradigms: aggregating preferences (e.g. voting, mechanism design) and aggregating information (e.g. discussion, consulting, forecasting). Almost all formally-studied group decisionmaking mechanisms fall under one paradigm or the other, but not both. We consider a public projects problem with the objective of maximizing utilitarian social welfare. Decisionmakers have both preferences, modeled as utility functions over the alternatives; and information, modeled as Bayesian signals relevant to the alternatives’ external welfare impact. Aligning incentives is highly challenging because, on the one hand, agents can provide bad information in order to manipulate the mechanism into satisfying their preferences; and on the other hand, they can misreport their preferences to favor selection of an alternative for which their information rewards are high.
We propose a two-stage mechanism for this problem. The forecasting stage aggregates information using either a wagering mechanism or a prediction market (the mechanism is modular and compatible with both). The voting stage aggregates preferences, together with the forecasts from the previous stage, and selects an alternative by leveraging the recently-studied Quadratic Transfers Mechanism. We show that, when carefully combined, the entire two-stage mechanism is robust to manipulation of all forms, and under weak assumptions, satisfies Price of Anarchy guarantees. In the case of two alternatives, the Price of Anarchy tends to $1$ as natural measures of the “size” of the population grow large. In most cases, the mechanisms achieve a balanced budget as well. We also give the first nonasymptotic Price of Anarchy guarantee for the Quadratic Transfers Mechanism, a result of independent interest.
@@ 1 Introduction
We study the public projects problem in which a group of decisionmakers is faced with selecting one of multiple alternative projects to undertake. Classically, the objective is to aggregate the preferences of the decisionmakers in order to make a choice which most benefits the group as a whole. Here, preferences will be modeled as quasilinear utility functions in a standard welfare-maximization framework, where a mechanism such as VCG might be employed.
@@
Instead of preferences, a different philosophy for group decisionmaking is to focus on aggregating information about which alternative is the best. For example, a committee could make a decision either by voting (preferences), or by discussing until consensus is reached (information). In the information-aggregation paradigm, decision markets (Othman and Sandholm, 2010) have been proposed as a group decisionmaking mechanism. There, forecasts about the impact of each alternative are produced by running prediction markets – financial markets whose prices reflect predictions – and then a decision is automatically chosen based on the predicted-best alternative. We likewise focus on predictions as a source of credible, verifiable information.
@@
A public projects problem can generally benefit from aggregation of both preferences and information. Can we design a formal mechanism for groups to make decisions based on both? There are only a few formal proposals in this spirit (see Section 1.2). When agents can serve as both information aggregators and preference-driven decisionmakers, they may have incentives to manipulate in both roles.
Model and main result.
In this paper, we will propose a mechanism that aggregates both preferences and predictions to select a public project. We consider monetary mechanisms and assume quasilinear utility, seeking social welfare guarantees. In our setting, each participant $i$ has a value $v_{k}^{i}\geq 0$ for each project $k$, with $V_{k}=\sum_{i}v_{k}^{i}$. Meanwhile, project $k$ has some external welfare impact $B_{k}$ representing the impact on the social welfare of non-decisionmakers. For example, $B_{k}$ could reflect an externality imposed by project $k$ such as carbon emissions or another environmental impact (see Section 1.1). In addition to their private values, the participants have private information, modeled as signals, relevant to each $B_{k}$. We consider the expected external impact of alternative $k$ conditioned on all signals as $B_{k}^{*}$. We seek to aggregate the preferences $\{v_{k}^{i}\}$ and expected external impacts $\{B_{k}^{*}\}$ in order to maximize total social welfare: $\arg\max_{k}(V_{k}+B_{k}^{*})$, the sum of welfare of the participants and the expected external impact.
The Synthetic players QUAdratic transfer mechanism with Predictions (SQUAP) begins with an information aggregation phase, which may be a wagering mechanism or a prediction market, to elicit $B_{k}^{*}$ for each $k$. Then, it uses a “voting” stage, based on the recently-introduced Quadratic Transfers Mechanism (QTM) discussed below, to elicit preferences and select an alternative. We obtain the following result on the Price of Anarchy (the worst-case ratio between social welfare achieved by (1) an equilibrium of the mechanism and (2) the best alternative):
@@
In SQUAP, for the correct choice of parameters, for $m=2$ alternatives,
@@
More formally, $T=\frac{\max_{k}(V_{k}+B_{k}^{*})}{\max_{i,k}v_{k}^{i}}$, the ratio of first-best welfare to the largest individual value. Thus, as the total welfare grows large relative to the maximum influence of any one person’s preference, the Price of Anarchy tends to one. For the case of $m\geq 3$ alternatives, we also obtain general results for SQUAP as a function of the Quadratic Transfer Mechanism’s Price of Anarchy, which is an open problem to fully analyze. Eguia et al. (2023) obtain a Price of Anarchy asymptotically tending to $1$ for QTM for $m\geq 3$ under some assumptions, which implies via our results the same guarantee for SQUAP.
SQUAP does require assumptions on the mechanism’s knowledge in order to implement it. We give several suggestions for relaxing these assumptions; see Section 1.1. More broadly, the result proves that aligning incentives to base a decision on both preferences and predictions is possible, even when strategic participants act as both forecasters and decisionmakers.
1.1 Outline of Techniques
@@
Recent work (Eguia et al., 2019, 2023) has introduced and studied the Quadratic Transfers Mechanism (QTM) for public projects. Inspired by quadratic voting (Lalley and Weyl, 2018; Goeree and Zhang, 2017), the mechanism collects bids (“votes”) on each alternative, with cost rising quadratically, and picks an alternative according to a probabilistic “softmax” of the vote totals. By redistributing payments, the mechanism is budget-balanced.
@@
QTM turns out to have several nice properties for our two-stage problem that will be discussed below. Our first technical result is an analysis of the QTM including Price of Anarchy guarantees (Section 3). Our approach builds on the work of Eguia et al. (2019), which gives asymptotic welfare bounds in a model with agents drawn i.i.d. from bounded value distributions, where realized values are common knowledge. We consider any worst-case set of agents with no bounds or distributional assumptions, and we analyze pure-strategy Nash equilibrium.
Incorporating information on external welfare.
In Section 4, we take one step toward incorporating information into the QTM. Before discussing these results, we digress briefly to explain the model of information in this paper.
Broadly, information can be relevant to the decisionmakers’ own utilities and/or to an “external” welfare impact on non-decisionmakers. Consider shareholders of a non-profit corporation voting on a proposal that includes a significant climate impact. Shareholders’ utilities may incorporate both the revenue generated and the relevance of the climate impact to themselves. But in addition, the climate impact has an externality on non-decisionmakers. The charter of the non-profit could include an obligation to incorporate such externalities into decisions. The total welfare of alternative $k$ is therefore $V_{k}+B_{k}^{*}$, the welfare of the decisionmakers plus the expected external welfare impact.
We consider a standard Bayesian model of information where decisionmakers observe private signals (the full model is presented in Section 2). For simplicity in this paper, we consider information that is only relevant to $B_{k}$, the external welfare impact. We assume that decisionmakers fully know their own values before receiving any signals. This model still captures the inherent strategic tension between satisfying preferences and acting on good information. Further, our mechanism can easily be applied to a setting where decisionmakers’ own preferences depend on private information, such as in common-value settings. However, analyzing the strategic properties in such cases is left as an open problem.
Coming back to the results, in Section 4, we modify the QTM to incorporate external welfare impacts in a case where the expectations $\{B_{k}^{*}\}$ of these impacts are known public information and non-manipulable. We show that, by simulating synthetic agents, we can retain QTM’s welfare guarantees as measured against total welfare $\max_{k}(V_{k}+B_{k}^{*})$. Because the mechanism simulates participants in the mechanism, it requires some information about values in order to play correctly, but this information is relatively minimal.
@@
In Section 5 we define a two stage mechanism, Synthetic players QUAdratic transfer mechanism with Predictions (SQUAP). The first stage has agents participate in an information-aggregation mechanism, such as a wagering mechanism or prediction market, to produce estimates $\{\hat{B}_{k}\}$ of the external welfare impacts of the projects. To produce the final decision, the second stage runs the Synthetic Players QTM from Section 4 with the aggregated estimations from the first stage as input.
To prove a Price of Anarchy bound, we take the following steps. First, we consider two commonly-studied mechanisms for eliciting predictions: wagering mechanisms and prediction markets. In each, the penalty for misreporting predictions can scale unboundedly as the mechanism grows “large,” even for misreports with small impact on the welfare estimates. This ensures that decisionmakers, if their utilities are small relative to total social welfare, cannot gain more than $\epsilon$ by manipulating the final predictions, and even when they do, the final predictions are still highly accurate. Second, we use an importance-weighting technique from the decision markets literature (Chen et al., 2011) so that forecasters’ expected payoffs are independent of the decision of the mechanism. This technique ensures that manipulating votes (preferences) cannot influence the expected rewards from providing information. This property relies on the QTM because the mechanism always produces positive probabilities for each alternative. Under some mild assumptions, we provide a Price of Anarchy guarantee, our main result, based both on the spread of the values $T$ and the size of the information-aggregation mechanism. The latter can be tuned to match $T$, and it can be subsidized (to an extent) by the revenue from the QTM.
One point to emphasize is that both stages of our mechanism are non-truthful in equilibrium. However, we show that the information stage is arbitrarily close to truthful and forecasts are highly accurate, while the QTM’s welfare properties hold in non-truthful equilibrium.
@@ 1.2 Related Work
There is naturally a significant amount of work on public projects from a mechanism-design perspective. We focus on the setting of this paper, which is social welfare maximization with quasilinear-utility agents. A standard solution is to use the Vickrey-Clarke-Groves mechanism (VCG) (Vickrey, 1961; Clarke, 1971; Groves, 1973) which has an equilibrium that is truthful and maximizes social welfare. However, VCG has some undesirable properties, including lack of a Price of Anarchy bound (there exist arbitrarily bad equilibria) and uncertain revenue (it may be very large or zero). Literature addressing the problem with more nuance than our basic model considers the excludable and non-excludable cases, e.g. Ohseto (2000). The smoothness framework (Roughgarden et al., 2017) has been used to analyze Price of Anarchy of public projects, particularly in a combinatorial setting. In particular, Lucier et al. (2013) show that a first-price mechanism can achieve a Price of Anarchy of $1$ (i.e. guarantee optimal welfare) in a sequential setting where bidders iteratively submit public bids.
@@
Recent research has considered transfer-based mechanisms which use quadratic payments  (Weyl, 2013; Lalley and Weyl, 2018; Goeree and Zhang, 2017). Lalley and Weyl (2018, 2019) propose a binary-outcome quadratic voting mechanism and show it is asymptotically efficient. Weyl (2017) then study this mechanism’s sensitivity to collusion, fraud, and voter mistakes when the number of agents is large. A number of other works study the behavior or empirical performance of quadratic voting (Chandar and Weyl, 2019; Casella and Sanchez, 2019; Quarfoot et al., 2017; Goeree and Zhang, 2017).
Building on Lalley and Weyl (2019), the QTM in particular is proposed by Eguia et al. (2019) for $m\geq 2$ alternatives. The authors consider a setting where participants’ values for the $m$ alternatives are drawn i.i.d. from a distribution over $[0,1]^{m}$, and realized values are common knowledge. They prove that for an undetermined choice of the mechanism’s scale parameter $c$, equilibria exist, and show that the probability of selecting the highest-welfare alternative approaches 1 as the number of agents diverges. Eguia et al. (2023) then study the same mechanism in a similar setting but under Bayes-Nash mixed-strategy equilibrium, with similar results. The part of our work on the QTM builds heavily on the approaches in Eguia et al. (2019, 2023), but our results address the nonasymptotic regime without distributional assumptions. This includes concrete bounds with small constants.
Information and decisionmaking.
We first note that some classic problems can be viewed as melding information with decisionmaking. In voting settings, jury theorems and related work (see e.g. Conitzer and Sandholm (2005)) interpret voting rules as aggregating information. A common-value auction (see e.g. Kremer (2002)) is a case of private information relevant to utilities and decisionmaking as well.
A number of works study mechanisms for eliciting information from experts and using it to make a decision. The most relevant mechanisms for this paper are decision markets (Hanson, 1999; Othman and Sandholm, 2010), which simultaneously operate a prediction market for each alternative $k$. (A prediction market is a financial market designed to aggregate beliefs into a consensus forecast, reflected in the prices of the financial products being traded (Hanson, 2003).) The decision market mechanism selects the alternative that is best according to the aggregated predictions (or uses a similar rule). Even without preferences over the alternatives, decision markets can have complex incentive misalignment problems, although these can be fixed by the importance-weighting technique of Chen et al. (2011).
@@
There are many works in which a single decisionmaker elicits information from one or more experts prior to making a decision, and the experts wish to influence the decision. This includes the “cheap talk” model of Crawford and Sobel (1982) as well as Bayesian Persuasion (Kamenica and Gentzkow, 2011). Somewhat closer to our setting, Oesterheld and Conitzer (2020) consider “decision scoring rules” that incentivize truthful predictions and recommendations by an expert.
For aggregating information from a group, e.g. Gerardi et al. (2009) consider a Bayesian setting in which a decisionmaker elicits information from a number of experts and aggregates it to make a decision. To incentivize truthfulness, the decisionmaker with small probability audits an expert’s report against the others (similar to peer prediction (Miller et al., 2005)) and picks an alternative desired by the audited agent. Such settings typically do not face the concern in this paper, where the decisionmakers are a large public group and may overlap with the forecasters.
@@
A few works are much closer to the spirit of our motivation, if technically quite different. They generally analyze voter preferences as a function of the information to be aggregated. In contrast, we model preferences as fixed with information as an orthogonal axis. In the “Wisdom-of-the-Crowd Voting Mechanism” of Schoenebeck and Tao (2021), each participant is both a voter and a holder of private information. The mechanism elicits both information and a preference in a single shot, then aggregates both to make a binary decision. The authors show that, with high probability, the mechanism selects despite strategic behavior the “majority wish” alternative that more voters would prefer if fully informed.
@@
Amanatidis et al. (2022) consider a similar problem motivated by blockchain applications. Voters have private information and participate in an approval vote among $m\geq 2$ alternatives. As with Schoenebeck and Tao (2021), information aggregation takes place within the mechanism rather than through an explicit phase prior to voting. The authors prove that the mechanism can achieve a Price of Anarchy of $\frac{1}{2}$.
@@
In Jackson and Tan (2013), a majority-vote between two alternatives takes place after a “deliberation” stage. The experts who take part in deliberation cannot also be voters (unlike in our model), and cannot misreport arbitrarily, but only choose to either reveal or hide their information in order to influence the voters. Other works in this spirit are Alonso and Câmara (2016); Schnakenberg (2015).
2 General Setting and Mechanism
We consider the problem of public projects with predictions, in which a group of $n$ agents needs to select one of $m\geq 2$ alternatives.
We will use $i,j$ to denote generic agents and $k,l$ to denote generic alternatives. Let $v_{k}^{i}\in\operatorname{\mathbb{R}}_{\geq 0}$ be agent $i$’s value for alternative $k$. We assume that all values are nonnegative. We consider a Nash equilibrium setting where values are common knowledge among the agents. Denote $\mathbf{v}^{i}=(v_{1}^{i},v_{2}^{i},\ldots,v_{m}^{i})$ as the vector of agent $i$’s values across all $m$ alternatives. The aggregate value for alternative $k$ is $V_{k}:=\sum_{i=1}^{n}v_{k}^{i}$.
Each agent also receives a signal $S_{i}$ before the mechanism begins. Let random variable $B_{k}\in\operatorname{\mathbb{R}}_{\geq 0}$ represent the impact to external welfare conditional on the $k$th alternative being chosen. Signals are drawn jointly from a common-knowledge prior $\mathcal{D}$ jointly along with the welfare impacts, i.e. we have $(S_{1},\dots,S_{n},B_{1},\dots,B_{m})\sim\mathcal{D}$. We let $\mathcal{D}_{k}$ be the marginal distribution of $\mathcal{D}$ over alternative $B_{k}$. The mechanism does not know $\mathcal{D}$.
The expected external welfare impact of alternative $k$, given all available information, is $B_{k}^{*}:=\operatorname{\mathbb{E}}[B_{k}\mid S_{1},\dots,S_{n}]$. The expected social welfare of alternative $k$ is $W_{k}:=V_{k}+B_{k}^{*}$, a random variable depending on the signals. The objective is to maximize expected social welfare (referred to as simply social welfare). We assume WLOG that $W_{1}\geq\cdots\geq W_{m}$, so that choosing alternative $1$ maximizes social welfare.
Our mechanism: framework.
We will consider mechanisms with two stages:
  1. 1.
    Forecasting. In the first stage, agents participate in an aggregation mechanism. We will formalize the mechanisms we use in Section 5. The aggregation mechanism produces a vector of aggregated forecasts, the external welfare estimates $\mathbf{\hat{B}}=(\hat{B}_{1},\dots,\hat{B}_{m})$. The aggregation mechanisms we consider, prediction markets and wagering mechanisms, reward participants based on the eventual observation of the actual outcome [^1] In reality, the welfare $B_{k}$ may not be directly observable, but we may use estimates or proxies for $B_{k}$ instead and modify the mechanism accordingly. $B_{k}$ once an alternative $k$ is chosen. At that time, the mechanism collects a payment $\pi^{i,1}$ from each agent $i$ as a function of the reports to the mechanism, the eventually-selected alternative $k$, and the realized outcome of $B_{k}$.
  2. 2.
    Decisionmaking. In the second stage, agents participate in the decisionmaking mechanism. For this stage, we will use a variant of the Quadratic Transfers Mechanism, defined in Section 3. The decisionmaking mechanism, which uses the external welfare estimates $\{\hat{B}_{k}\}$ as part of its process, outputs a probability distribution $\mathbf{p}\in\Delta_{m}$. It also collects a payment $\pi^{i,2}$ from each agent $i$ as a function of the reports to the mechanism. The final group decision is $k$ drawn randomly according to $\mathbf{p}$.
If the joint mechanism selects alternative $k$ under $\mathbf{p}$, the social welfare of the mechanism is $\operatorname{\mathbb{E}}_{k\sim\mathbf{p}}W_{k}:=\operatorname{\mathbb{E}}_{k\sim\mathbf{p}}[V_{k}+B_{k}^{*}]$. Each agent’s ex-post utility when alternative $k$ is selected is $v^{i}_{k}-\pi^{i,1}-\pi^{i,2}$.
Approach: backward induction on the stages.
As discussed, there are conceptual incentive challenges in designing a mechanism for this context. By using the two-stage approach, we modularize the problem to an extent. In Section 5, we will be able to use properties of wagering mechanisms and prediction markets to bound the impact of manipulation in stage 1. However, before considering stage 1, we need to solve the second stage of the game. Thus, it is useful for us to begin with analyzing the Quadratic Transfers Mechanism in Section 3. Then we will extend it in Section 4 to incorporate external welfare estimates (from non-strategic sources).